Task 1 – Commercial banks in United Economy have total deposits of AED 300 billion. Their reserves are AED 15 billion, two- thirds of which are with the Central Bank as deposits. There are AED 30 billion notes outside the banks. There are no coins!

Calculate-
a) The monetary base.
b) The banks’ reserve ratio.
c) Currency drain as a % of deposits.
1
Expert's answer
2016-04-13T08:09:05-0400
a) Monetary Base = Currency + Bank Reserves
Monetary base = 30+ 15 = 45 billions

b) Bank reserve ratio = Bank reserves / Deposits
Bank reserve ratio = 15/300 = 0.05

c) Currency drain as a % of deposits = Currency·100%/Deposits
Currency drain as a % of deposits = 30 · 100/300=10%